### v1

# Water Detection in the Interstellar Object 3I/ATLAS

Abstract
We report the first detection of water activity in the third confirmed interstellar object, 3I/ATLAS, based on ultraviolet imaging with the Neil Gehrels-Swift Observatory. Observations acquired with the Ultraviolet/Optical Telescope on 2025 July 31st - Aug 1st revealed OH (A — X) emission near 3085 Å. The water production rate results highly depend on the reddening assumption. For a reddening of 38.6% between 5437.8 Å and 3325.7 Å, the water production rate is molecules s-1 (40 kg s-1) at a heliocentric distance of 3.51 au. This places 3I/ATLAS among the few comets with confirmed OH emission beyond 3 au, where water ice sublimation is typically inefficient. The inferred production rate is consistent with an active area of at least 19 km2, assuming equilibrium sublimation. Based on current upper limits of the nucleus’ radius, this requires that over 20% of the surface is active, which is larger than activity levels observed in most solar system comets. Contemporaneous near-infrared spectroscopy indicates the presence of large icy grains in the coma, which may serve as an extended source of water vapor. The detection of OH emission prior to any CN detection is unusual and may reflect differences in grain-driven outgassing or volatile inventory compared to typical comets. While similar behavior has been observed in solar system comets, the mechanisms controlling distant activity and the storage and release of volatiles remain poorly understood. If 3I/ATLAS’ coma continues to be dominated by H2O, supporting the early and low-metallicity formation hypothesis, the derived large size of the nucleus could be indicative of a key knowledge gap in low-metallicity system planetesimal formation and loss mechanisms.
1 Introduction
The discovery of the third interstellar object, 3I/ATLAS, on 1 July 2025 initiated a broad characterization campaign across the globe (L. Denneau, 2025; D. Z. Seligman et al., 2025). Following the lessons learned from the prior interstellar objects 1I/‘Oumuamua and 2I/Borisov observing campaigns were initiated to rapidly capture its initial brightness, morphology, lightcurves, color, and optical and near-infrared spectrum (‘Oumuamua Team, 2019; A. Fitzsimmons et al., 2023). Given the apparent brightness and early extension of the coma the production of gas was assumed and searched for (A. Alvarez-Candal et al., 2025; C. Opitom et al., 2025) but not found. Characterizing the early activity of interstellar objects is essential for understanding their chemical and physical evolution during solar approach, as it possibly represents the first time the object has been significantly heated in their very long dynamic lifetimes (M. J. Hopkins et al., 2025; A. G. Taylor & D. Z. Seligman, 2025).
Inbound activity for our own Oort Cloud solar system comets is often more predictable. Hypervolatiles like CO and CO2 drive activity beyond the water ice line (T. Ootsubo et al., 2012; O. H. Pinto et al., 2022), which can entrain small icy grains into the coma that sublimate to produce H2O (M. F. A’Hearn et al., 2011). As comets approach the Sun, water increasingly dominates the observed gas composition. This is not necessarily true for all solar system comets (see C/1908 R1 (Morehouse), C/2016 R2 (PANSTARRS), and C/2009 P1 (Garradd); (A. J. McKay et al., 2019; N. Biver et al., 2018; D. Bodewits et al., 2014), nor was it the case for interstellar comet 2I/Borisov, which transitioned from a equally mixed H2O/CO coma pre-perihelion to a CO-dominated coma post-perihelion (D. Bodewits et al., 2020; Z. Xing et al., 2020; M. A. Cordiner et al., 2020). The change in composition was only captured thanks to the fortunate timing of UV observations by HST and the Neil Gehrels-Swift Observatory that monitored CO and H2O production. A similar monitoring campaign was planned for the next interstellar object to establish a chemical baseline for comparison.
The Ultraviolet and Optical Telescope (UVOT) onboard the Neil Gehrels-Swift Observatory has proven to be an excellent instrument for monitoring the OH (A – X) band at 3085 Å (K. Mason et al., 2007; J. A. Carter et al., 2012; D. Bodewits et al., 2014; Z. Xing et al., 2020; D. Bodewits et al., 2023). Despite its modest aperture, UVOT’s sensitivity is enhanced by its position above Earth’s atmosphere, free from telluric absorption and sky background, enabling deep integrations and effective stacking. Because it is flown above the atmospheric extinction, UVOT has the equivalent sensitivity of a 4-m telescope on the ground111https://swift.gsfc.nasa.gov/proposals/tech_appd/swiftta_v12.pdf. UVOT was used to derive the H2O production rate of 2I/Borisov during its flyby of the solar system in 2019/2020, and it was the dual measurements of UVOT and HST COS that showed the duality of Borisov’s chemical composition around perihelion (D. Bodewits et al., 2020; Z. Xing et al., 2020).
Here, we present our initial UVOT observations of 3I/ATLAS taken on July 31 and August 1, 2025 that show the onset of water activity in the third interstellar object. In the following sections we describe the observations and analysis, place our results in context with the current 3I/ATLAS literature, and compare and contrast with both the previous interstellar objects and Oort Cloud comet populations.
2 Observation and Analysis
We observed 3I/ATLAS with the Ultraviolet/Optical Telescope (UVOT; P. W. Roming et al., 2000) on board the Neil Gehrels-Swift Observatory (N. Gehrels et al., 2004). UVOT is a 30-cm Ritchey–Chrétien telescope with a 17 arcmin17 arcmin field of view and a plate scale of 0.502 arcsec/pixel, although most images are taken in ‘image mode’ with 22 on-board binning, resulting in an effective plate scale of 1.004 arcsec/pixel. It is equipped with 11 broadband filters and two grisms, covering wavelengths from 1600 to 8000 Å.
Observations were obtained between 2025 July 31 05:24 UTC and 2025 August 01 09:49 UTC, when the comet was at a heliocentric distance of 3.51 au and a geocentric distance of 2.82 au (Table 1). We used UVOT’s UVW1 filter (central wavelength = 2600 Å, FWHM = 693 Å), which includes most of the OH (A - X) emission band centered near 3085 Å. To estimate and subtract the continuum contribution within the UVW1 bandpass, we also acquired images using the V-band filter ( = 5468 Å, FWHM = 769 Å).
Swift does not track moving objects, so comets are typically observed with brief 200-s exposures. During our observations, 3I had a relatively high apparent motion of 1.79 to 1.86 arcsec/min, corresponding to a motion of about 6 arcsec in a typical 200-s exposure and thus exceeding UVOT’s point spread function (PSF; 2.18 arcsec FWHM in V and 2.37 arcsec FWHM in UVW1 according to A. A. Breeveld et al. (2010)). To avoid blurring and further dilution of the surface brightness, we acquired the UVW1 images in ‘event-mode’, where the arrival time of every photon is time tagged. Due to the telemetry load constraints of Swift and the high brightness in the V band images, we could not use event mode for the V band observations and instead acquired short-exposure images in the ‘image mode’.
| Mid Time (UTC) | (AU) | (km s-1) | (AU) | S–T–O (deg) | (days) | Filter | Images | Exp. Time (s) | ||||
[…]
To mitigate motion blur, we divided photons in each event-mode UVW1 image into 30-second time slices, limiting motion within each slice to less than 1 arcsec. We then aligned all slices by the target’s nucleus position (from JPL/Horizons) and summed them to produce motion-corrected UVW1 images. These UVW1 images have a pixel scale of 0.502 arcsec pixel-1, and we binned them by 2×2 to match the 1.004 arcsec pixel-1 scale of V-band images obtained in the image mode for subsequent image subtraction. For V-band images, such motion correction was not possible without event-mode images.
Next, we discarded all images in which the target was contaminated by background stars, then aligned the remaining images on the nucleus’ position and stacked them separately for each filter to improve the SNR ratio. The resulting UVW1 and V-band stacked images, shown in Fig. 1, have equivalent exposure times of 9055 s and 2713 s, respectively. Both images show clear detections of 3I/ATLAS. In the V-band image, the comet appears elongated along the direction of motion, as expected. The UVW1 image, free from smearing, shows a spatially extended source with a FWHM of 5 arcseconds and a slight asymmetry across the comet–Sun axis.
In the stacked images, we noticed that the target positions deviated from JPL/Horizons predictions. This offset, also present for stars in raw images, likely stems from a small error in the astrometric solution produced by the UVOT pipeline. The UVW1 stacked image showed particularly large errors (approximately 5 arcsec), suggesting greater uncertainty in event mode. We fitted the target with a 2D Gaussian to locate the nucleus position for further analyses.
To identify stars around the target, we applied a 3-sigma detection threshold requiring at least 4 connected pixels. The background was measured using a star-masked annulus with an inner radius of 30 arcsec and an outer radius of 40 arcsec (corresponding to 60,000 - 80,000 km). The derived backgrounds are counts s-1 pixel-1 for the V-band and counts s-1 pixel-1 for UVW1. After background subtraction, aperture photometry with a 10-arcsec radius yielded counts s-1 for UVW1, while an elongated 10-arcsec aperture accounting for motion blur yielded counts s-1 for V-band.
We converted the V-band net count rate to flux and magnitude using Swift’s effective area (corrected for UVOT’s ongoing sensitivity decline; A. A. Breeveld et al. (2011)) and a target spectrum model. We adopted a linearly reddened solar spectrum to model the target spectrum with the default solar spectrum included in the sbpy package (M. Mommert et al., 2019), and used a radius of 15 cm to derive the UVOT collecting area. The calculated flux, together with the same target spectrum model and Vega spectrum (R. C. Bohlin, 2014), was used to derive the apparent Vega magnitude in the V band of UVOT. For the V filter, the derived V-band fluxes and magnitudes are nearly insensitive to different reddening assumptions, with result variations less than 0.4%. Therefore, the uncertainty introduced by the reddening assumption is negligible for the photometry. The measured V-band flux is erg s-1 cm-2, corresponding to a V magnitude of mag.
The net count rate of the V image also enables an estimate of the reflected continuum contribution to the UVW1 flux, allowing us to isolate the OH (A – X) emission:
where is the count rate attributable to OH, and are the count rates measured in the UVW1 and V images, respectively, as described above, and is the solar continuum count rate ratio between the filters. The derived OH count rate depends heavily on the ratio factor , which is a function of the reddening per 1000 Å between the photon-weighted effective wavelengths of V and UVW1 filters. These effective wavelengths, calculated using the sensitivity-corrected effective area and a solar spectrum, are 5437.8 Å for V and 3325.7 Å for UVW1. To determine the count rate ratio , we convolved a solar spectrum linearly reddened by with the respective effective areas.
The OH count rates are converted to the emitted flux using a hydroxyl spectral model (D. Bodewits et al., 2019), then to luminosity using the geocentric distance. We subsequently derived the total number of OH molecules within the 10-arcsec radius aperture using fluorescence efficiencies (D. G. Schleicher & M. F. A’Hearn, 1988) appropriate for the specific geocentric velocities, with a scaling applied for the heliocentric distance.
To derive water production rates from the total number of OH molecules in the aperture, we applied the vectorial model with a random OH ejection kick of 1.05 km s-1 by taking photodissociation of H2O into account (M. C. Festou, 1981). We adopted a H2O outflow velocity of km s-1, photodissociation lifetimes of 86 000 s for H2O and 129 000 s for OH, and a branching ratio of 0.93 for water dissociation into OH (M. R. Combi et al., 2004). We compared observed and model-predicted OH molecular numbers within the same aperture. The observed water production rate was then obtained by scaling the modeled water production rate with this ratio.
3 Results and Discussion
The UVOT-V magnitude of mag corresponds to a Johnson V-mag of . This results in an of 174 cm after correcting for the phase angle (M. F. Ahearn et al., 1984; D. G. Schleicher & A. N. Bair, 2011). If we apply the V-R color of 0.51 mag measured on July 23 from T. Santana-Ros et al. (2025), the is 180 cm in the R band. The discrepancy between our results and the 300 cm reported by T. Santana-Ros et al. (2025) could originate in temporal coma development or aperture size differences.
Our water production rate measurement depends strongly on the assumed reddening between the effective wavelengths of the V and UVW1 filters (5437.8 Å and 3325.7 Å, respectively), as shown in the left panel of Fig. 2. The error bars represent 1- stochastic uncertainties. A possible water detection emerges for reddening assumptions larger than , while a clear 3- detection is achieved for reddening larger than .
We then investigated reddening measurements of 3I/ATLAS from other telescopes, presented in the middle panel of Fig. 2. The x-axis shows reddening values (percent per 1000 Å), while the y-axis indicates the central wavelength of the measurement bands. Different marker shapes denote literature sources, with colors representing the comet’s heliocentric distance during the observations (colorbar on the far right). A. Alvarez-Candal et al. (2025) measured per 1000 Å across 3200–5600 Å, with a central wavelength closest to ours (4381.8 Å). We show this reddening measurement in the left panel with a green dashed line. Other reddening measurements typically range from 10% to 25%, with higher reddening values at shorter wavelengths.
This trend of increasing reddening toward shorter wavelengths is unsurprising. For example, a trend of increased reddening at shorter central wavelengths is presented in D. Jewitt & K. J. Meech (1986). Directly measured values in the UV band corroborate this trend, with values of 60–80%/1000 Å near 2950 Å from C/1983 O1 (Černis) (P. D. Feldman & M. F. A’Hearn, 1985) and 70%/1000 Å near 2950 Å from 1P/Halley (P. D. Feldman et al., 1987). In contrast, ground-based measurements in the B-V domain hover around 20% depending on cometary activity, size, and composition (C. Opitom et al., 2025; A. D. Storrs et al., 1992; D. Jewitt & K. J. Meech, 1986).
This trend has two sources. First, reddening is defined as the reflectance spectral slope normalized by the mean reflectance, not the raw slope. This normalization causes wavelength-dependent variations even for linear reflectance spectra, with higher values at shorter wavelengths. Second, reflectance spectra are often non-linear and they steepen at shorter wavelengths, further increasing reddening values. For 3I/ATLAS, reflectance spectra in several studies also exhibit such steepening at short wavelengths (c.f. T. Santana-Ros et al., 2025; A. Alvarez-Candal et al., 2025).
To obtain reddening values at the V and UVW1 effective wavelengths for water production rate determination, we converted the middle panel’s reddening measurements from other wavelengths to UVOT wavelengths. For each measurement, we constructed a linear reflectance spectrum that, when multiplied by the solar spectrum, produces a red dust spectrum with the reported reddening value at that measurement wavelength. We then forward-modeled this reddened spectrum through Swift’s UVW1 and V filters and computed the reddening at UVOT effective wavelengths. Uncertainties in the converted reddening values were derived using error propagation.
We present the converted results in the right panel of Fig. 2. The y-axis still shows the original measurement wavelengths, while the x-axis displays the reddening converted to UVOT wavelengths. All reddening values increased after conversion, as expected from the reddening definition. Nearly all converted values exceed 20%, with a mean of 30.3%. At the 20% lower limit, the water production rate is molecules/s (SNR = 2.1), while at the mean reddening value, it reaches with SNR = 3.8, indicating clear detection. Our calculations assume linear reflectance spectra. When considering that reflectance spectra will steepen towards shorter wavelengths, actual reddening values would be higher, strengthening the water detection. The conversion accuracy depends on the wavelength difference between the original measurement (y-axis) and UVOT’s middle effective wavelength (4381.8 Å) — smaller differences yield higher accuracy. The measurement by A. Alvarez-Candal et al. (2025), nearest to our UVOT filter set (3200-5600 Å), converts to 38.6%, corresponding to a definitive detection of with SNR = 5.0, which corresponds to 40 kg s-1 at a heliocentric distance of 3.51 au.
The ratio log[Af/Q(H2O)] = -24.9 places 3I/ATLAS among the most dust-rich comets observed, comparable to a small subset that includes C/1995 O1 (Hale-Bopp), C/1991 B1 (Shoemaker-Levy), and C/2009 P1 (Garradd) (M. F. A’Hearn et al., 1995; T. Farnham et al., 1997; Y. R. Fernandez, 2000; D. Bodewits et al., 2014). We note, however, that the Af values in the M. F. A’Hearn et al. (1995) sample were not corrected for phase angle and OH production rates were used there, which will be lower than water production rates.
Using a sublimation model222https://pdssbn.astro.umd.edu/tools/ma-evap/index.shtml (J. J. Cowan & M. F. A’Hearn, 1979) and assuming constant solar elevation (as would occur for slow rotation or a Sun-facing pole), a Bond albedo of 0.05%, and unit infrared emissivity, we estimate the minimum required active area to be 19.47 km2, equivalent to a circular area with radius 2.47 km.
Observations acquired with the Hubble Space Telescope placed an upper limit on the nucleus’ radius of 2.8 km (D. Jewitt et al., 2025). If the water is produced directly by sublimation of ice within the nucleus, it would require that 20% of the surface contributes, whereas most solar system comets have active fractions around 3-5% (M. F. A’Hearn et al., 1995; H. U. Keller et al., 2015).
Gas detections beyond 2.5 au are uncommon due to the inefficiency of water sublimation at those distances. Only a few comets, such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), have shown OH detections at such large distances. OH emission from Bowell was detected at 5.3 au (M. F. Ahearn et al., 1984), requiring an implausibly large active area (1.6 107 km2) unless icy grains contributed significantly to water release. The OH production showed non-monotonic evolution, suggesting transient grain-driven activity. Similarly, Swift and SOHO/SWAN observations of comet Garradd revealed elevated water production beyond 3.4 au compared to narrow-slit spectroscopy, consistent with extended sources such as icy grains in the coma (M. R. Combi et al., 2013; D. Bodewits et al., 2014).
Near-infrared spectroscopic observations of 3I/ATLAS obtained with Gemini South/GMOS on July 5, 2025 UTC, and IRTF/SpeX on July 14, 2025 UTC, suggest the presence of large icy grains in the coma (B. Yang et al., 2025). If these grains are the primary source of the water and OH observed in the coma, they may explain why OH (A – X) emission was detected before any CN emission. Typically, CN is among the most distant fragment species detected in comets due to the high fluorescence efficiency of its B – X (0,0) violet system near 3883 Å, which also benefits from lower atmospheric extinction and higher detector sensitivity compared to the OH emission near 3085 Å (c.f. A. L. Cochran & W. D. Cochran, 1991). Large, dark icy grains in the coma are more efficiently heated than the nucleus, leading to preferential loss of their more volatile components. A similar process was observed at 67P/Churyumov–Gerasimenko, where the transport of material from the southern to the northern hemisphere resulted in enhanced H2O abundances relative to more volatile species such as CO2, CO, and HCN, implying that these more volatile ices were lost as lofted backfall material (H. U. Keller et al., 2017).
The physical processes that govern distant cometary activity and the retention or release of volatiles in such regimes remain poorly understood. Observational biases, limited statistics, and the challenge of detecting tenuous outgassing hinder progress. However, the unusual characteristics of 3I/ATLAS including early OH but no CN detection, strong coma reddening, and inferred large grains, highlight the need for long-term, multiwavelength monitoring. Simultaneous observations of gas and ice (e.g., via JWST) will be essential to disentangle the roles of native nucleus sublimation and extended grain activity, as well as probe the abundances of any hypervolatiles like CO and CO2. As noted by both M. J. Hopkins et al. (2025) and A. G. Taylor & D. Z. Seligman (2025), 3I/ATLAS may be water-rich, consistent with planetesimals formed in the outer regions of low-metallicity planetary systems. Our results offer one of the first direct tests of that hypothesis, finding indirect evidence for a large H2O production rate, but it remains to be seen if 3I/ALTAS’ coma is dominated by the other volatiles.
The presence of substantial OH, and the inferred production of H2O, can then be used in direct comparison to 2I/Borisov. Borisov had a much lower intercept speed with our solar system and likely came from a more recently formed system; the high CO/H2O ratio post-perihelion was seen as evidence of formation near the CO ice line, likely of an M-type star (D. Bodewits et al., 2020). 3I/ATLAS, which is moving nearly twice as fast as Borisov, is also likely to show asymmetric activity around perihelion as the thermal wave penetrates into deeper reservoirs. The post-perihelion behavior of 3I will expose more pristine materials, unaltered from the inbound journey. 3I/ATLAS has a H2O production rate at 3.5 au similar to the maximum value measured for 2I/Borisov at 2.4 au (J. Crovisier et al., 2019; Z. Xing et al., 2020), and therefore we propose the following tests of the H2O-rich hypothesis:
1.
If 3I/ATLAS is H2O-rich, indicating an older age and formation in a low-metallicity system, then the H2O production rate will peak near perihelion, and only trace amounts of high-metallicity volatiles like CO,CO2, CN and hydrocarbon species will be found.
2.
If 3I/ATLAS is hyper-volatile rich (CO/CO2), indicating that its dynamic age does not necessarily represent the metallicity of where it formed, the H2O production rate will drop precipitously postperihelion as more volatile species are sublimated from depth. This would be similar to the CO/H2O observations of 2I/Borisov (D. Bodewits et al., 2020).
If the first hypothesis is born out, and 3I/ATLAS is indeed from an older low-metallicity system, than this may suggest that there is a part of the planetesimal formation or removal process that may have ejected more mass from forming systems in order to properly place the large size of 3I into context (A. G. Taylor & D. Z. Seligman, 2025). Given the ISO production of low-metallicity systems would likely come from planetary interactions (S. N. Raymond et al., 2018) or stripping by stellar flybys (S. Pfalzner et al., 2021) an overabundance of large, H2O-rich ISO’s could serve as a key indicator of ejection conditions.
4 Conclusion
We report the detection of OH (A – X) emission in the coma of 3I/ATLAS using ultraviolet imaging with the Neil Gehrels-Swift Observatory. This confirms the presence of water in the third known interstellar object. The derived production rate of molecules s-1 (40 kg s-1) is notable given the object’s large heliocentric distance of 3.51 au, where water sublimation is typically inefficient. This level of activity is consistent with a reasonably-sized active area assuming equilibrium sublimation, but could also result from the presence of large icy grains in the coma, as suggested by contemporaneous near-infrared observations. If the coma of 3I/ATLAS remains H2O dominated and depleted in higher-metallicity volatiles like CO, CN, and CO2, the relatively large size of the interstellar object may be a telling clue about the planetesimal and mass loss processes of early low-metallicity systems.
The detection of OH emission prior to any detection of CN is unusual and may indicate differences in grain composition or volatile distribution compared to solar system comets (C. Opitom et al., 2025; A. Alvarez-Candal et al., 2025). While similar behavior has been observed in comets such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), the processes governing distant activity and the release of volatiles remain poorly constrained. Grain-driven activity, in particular, may play a more significant role in interstellar objects than previously assumed. This detection demonstrates the critical role of UV space-based observations in tracking the onset and evolution of activity in distant or dynamically unusual objects. Continued monitoring of 3I/ATLAS will help clarify the thermal and compositional properties of interstellar ices and improve our understanding of volatile retention and release in planetary systems beyond our own.

### v2

@@
-Water Detection in the Interstellar Object 3I/ATLAS
+Water Production Rates of the Interstellar Object 3I/ATLAS
 Abstract
-We report the first detection of water activity in the third confirmed interstellar object, 3I/ATLAS, based on ultraviolet imaging with the Neil Gehrels-Swift Observatory.
-Observations acquired with the Ultraviolet/Optical Telescope on 2025 July 31st - Aug 1st revealed OH (A — X) emission near 3085 Å.
-The water production rate results highly depend on the reddening assumption.
-For a reddening of 38.6% between 5437.8 Å and 3325.7 Å, the water production rate is molecules s-1 (40 kg s-1) at a heliocentric distance of 3.51 au.
-This places 3I/ATLAS among the few comets with confirmed OH emission beyond 3 au, where water ice sublimation is typically inefficient.
-The inferred production rate is consistent with an active area of at least 19 km2, assuming equilibrium sublimation.
-Based on current upper limits of the nucleus’ radius, this requires that over 20% of the surface is active, which is larger than activity levels observed in most solar system comets.
-Contemporaneous near-infrared spectroscopy indicates the presence of large icy grains in the coma, which may serve as an extended source of water vapor.
-The detection of OH emission prior to any CN detection is unusual and may reflect differences in grain-driven outgassing or volatile inventory compared to typical comets.
-While similar behavior has been observed in solar system comets, the mechanisms controlling distant activity and the storage and release of volatiles remain poorly understood.
-If 3I/ATLAS’ coma continues to be dominated by H2O, supporting the early and low-metallicity formation hypothesis, the derived large size of the nucleus could be indicative of a key knowledge gap in low-metallicity system planetesimal formation and loss mechanisms.
+We report the detection of water activity in the third confirmed interstellar object, 3I/ATLAS, based on ultraviolet imaging with the Neil Gehrels Swift Observatory’s Ultraviolet/Optical Telescope (UVOT).
+Assuming a reddening of 29% between 3325.7 Å and 5437.8 Å, measurements on 2025 July 31–August 1 yielded a first, marginal detection of OH (A – X) emission near 3085 Å, corresponding to a water production rate of molecules s-1.
+The subsequent visit on 2025 August 18–20 revealed a clear OH detection, implying a higher water production rate of molecules s-1 (40 kg ) at a heliocentric distance of 2.90 au.
+This places 3I/ATLAS among the few comets with confirmed OH emission beyond 2.5 au, where water ice sublimation from the nucleus is typically inefficient.
+The inferred production rate at 2.9 au implies an active area of at least 7.8 km2, assuming equilibrium sublimation.
+This requires that over 8% of the surface is active, which is larger than activity levels observed in most solar system comets.
+Contemporaneous near-infrared spectroscopy indicated the presence of icy grains in the coma, which may serve as an extended source of water vapor.
 1 Introduction
@@
 Seligman et al., 2025).
-Following the lessons learned from the prior interstellar objects 1I/‘Oumuamua and 2I/Borisov observing campaigns were initiated to rapidly capture its initial brightness, morphology, lightcurves, color, and optical and near-infrared spectrum (‘Oumuamua Team, 2019; A.
+Following the lessons learned from the prior interstellar objects 1I/‘Oumuamua and 2I/Borisov, observing campaigns were initiated to rapidly capture its initial brightness, morphology, lightcurves, color, and optical and near-infrared spectrum (‘Oumuamua Team, 2019; A.
 Fitzsimmons et al., 2023).
-Given the apparent brightness and early extension of the coma the production of gas was assumed and searched for (A.
-Alvarez-Candal et al., 2025; C.
+Given the apparent brightness and early extension of the coma the production of gas was assumed and searched for (C.
 Opitom et al., 2025) but not found.
@@
 Biver et al., 2018; D.
-Bodewits et al., 2014), nor was it the case for interstellar comet 2I/Borisov, which transitioned from a equally mixed H2O/CO coma pre-perihelion to a CO-dominated coma post-perihelion (D.
+Bodewits et al., 2014), nor was it the case for interstellar comet 2I/Borisov, which transitioned from an equally mixed H2O/CO coma pre-perihelion to a CO-dominated coma post-perihelion (D.
 Bodewits et al., 2020; Z.
@@
 Xing et al., 2020).
-Here, we present our initial UVOT observations of 3I/ATLAS taken on July 31 and August 1, 2025 that show the onset of water activity in the third interstellar object.
+Here, we present our initial UVOT observations of 3I/ATLAS taken on July 31 and August 19, 2025 that show the onset of water activity in the third interstellar object.
 In the following sections we describe the observations and analysis, place our results in context with the current 3I/ATLAS literature, and compare and contrast with both the previous interstellar objects and Oort Cloud comet populations.
@@
 It is equipped with 11 broadband filters and two grisms, covering wavelengths from 1600 to 8000 Å.
-Observations were obtained between 2025 July 31 05:24 UTC and 2025 August 01 09:49 UTC, when the comet was at a heliocentric distance of 3.51 au and a geocentric distance of 2.82 au (Table 1).
+We report here on observations conducted on two visits: 2025 July 31 – August 1, when the comet was at a heliocentric distance () of 3.51 au and a geocentric distance () of 2.82 au, and 2025 August 18–20, when it was at au and au.
+(Table 1).
 We used UVOT’s UVW1 filter (central wavelength = 2600 Å, FWHM = 693 Å), which includes most of the OH (A - X) emission band centered near 3085 Å.
@@
 Swift does not track moving objects, so comets are typically observed with brief 200-s exposures.
-During our observations, 3I had a relatively high apparent motion of 1.79 to 1.86 arcsec/min, corresponding to a motion of about 6 arcsec in a typical 200-s exposure and thus exceeding UVOT’s point spread function (PSF; 2.18 arcsec FWHM in V and 2.37 arcsec FWHM in UVW1 according to A.
+During our observations, 3I had a relatively high apparent motion of 1.74 to 1.85 arcsec/min, corresponding to a motion of about 6 arcsec in a typical 200-s exposure and thus exceeding UVOT’s point spread function (PSF; 2.18 arcsec FWHM in V and 2.37 arcsec FWHM in UVW1 according to A.
 A.
@@
 (2010)).
-To avoid blurring and further dilution of the surface brightness, we acquired the UVW1 images in ‘event-mode’, where the arrival time of every photon is time tagged.
-Due to the telemetry load constraints of Swift and the high brightness in the V band images, we could not use event mode for the V band observations and instead acquired short-exposure images in the ‘image mode’.
-| Mid Time (UTC) | (AU) | (km s-1) | (AU) | S–T–O (deg) | (days) | Filter | Images | Exp. Time (s) | ||||
-[…]
-[…]
+To avoid blurring and further dilution of the surface brightness, we acquired the UVW1 images in ‘event-mode’, where the arrival time of every photon is time tagged (T.
+S.
+Poole et al., 2008).
+Because of telemetry load constraints on Swift and the higher sky brightness in the V-band, we were unable to use event mode for the V-band observations and instead obtained short-exposure images in ‘image’ mode.
+| Visit | 07-31 | 08-19 |
+| Observational log | ||
+| Mid Time (UTC) | 2025-07-31 19:38:50 | 2025-08-19 07:52:11 |
+| (AU) | 3.51 | 2.90 |
+| (km/s) | ||
+| (AU) | 2.82 | 2.63 |
+| S–T–O angle (deg) | 13.65 | 20.33 |
+| UVW1 image numbera | 16 (15) | 14 (14) |
+| UVW1 exposure time (s)a | 9395 (8355) | 8206 (8206) |
+| V image numbera | 16 (13) | 14 (13) |
+| V exposure time (s)a | 3100 (2522) | 2708 (2517) |
+| Photometric Resultsb | ||
+| UVW1 net count rate (cnts/s) | ||
+| V net count rate (cnts/s) | ||
+| V flux (erg/s/cm2) | ||
+| V Magnitude | ||
+| V A(0)f (cm) | 202 | 337 |
+| OH Resultsc | ||
+| Number of OH (molecules) | ||
+| Q(H2O) (molec/s) | ||
+| Min. effective area (km2) | 10.6 | 7.8 |
+| Min. active fraction (%) | 11 | 8 |
+| Water loss mass (kg) | 22 | 40 |
+. Numbers in parentheses denote the equivalent number of images and exposure time used for analysis after excluding images contaminated by stars.
+. All photometric values in this table are measured within a 10-arcsec aperture; the V-band measurements used an elongated aperture of the same effective size to account for motion blur.
+. All OH-related results in this table are derived assuming a reddening of 29%.
 To mitigate motion blur, we divided photons in each event-mode UVW1 image into 30-second time slices, limiting motion within each slice to less than 1 arcsec.
@@
 For V-band images, such motion correction was not possible without event-mode images.
-Next, we discarded all images in which the target was contaminated by background stars, then aligned the remaining images on the nucleus’ position and stacked them separately for each filter to improve the SNR ratio.
-The resulting UVW1 and V-band stacked images, shown in Fig.
-1, have equivalent exposure times of 9055 s and 2713 s, respectively.
-Both images show clear detections of 3I/ATLAS.
-In the V-band image, the comet appears elongated along the direction of motion, as expected.
-The UVW1 image, free from smearing, shows a spatially extended source with a FWHM of 5 arcseconds and a slight asymmetry across the comet–Sun axis.
+Next, we discarded all images in which the target nucleus was contaminated by background stars.
+Each remaining image was divided into 50×50-pixel subregions, and stars located beyond 30 arcsec from the nucleus were identified in each subregion using a 3- threshold above the local background, requiring connected pixels.
+Each region was then expanded by 2 pixels to cover the PSF.
+Within 30 arcsec of the nucleus, contaminated areas were flagged manually.
+All identified regions were filled with the surrounding median.
+We aligned the remaining images on the nucleus’ position and stacked them separately for each filter to improve the SNR ratio.
 In the stacked images, we noticed that the target positions deviated from JPL/Horizons predictions.
 This offset, also present for stars in raw images, likely stems from a small error in the astrometric solution produced by the UVOT pipeline.
-The UVW1 stacked image showed particularly large errors (approximately 5 arcsec), suggesting greater uncertainty in event mode.
-We fitted the target with a 2D Gaussian to locate the nucleus position for further analyses.
-To identify stars around the target, we applied a 3-sigma detection threshold requiring at least 4 connected pixels.
-The background was measured using a star-masked annulus with an inner radius of 30 arcsec and an outer radius of 40 arcsec (corresponding to 60,000 - 80,000 km).
-The derived backgrounds are counts s-1 pixel-1 for the V-band and counts s-1 pixel-1 for UVW1.
-After background subtraction, aperture photometry with a 10-arcsec radius yielded counts s-1 for UVW1, while an elongated 10-arcsec aperture accounting for motion blur yielded counts s-1 for V-band.
-We converted the V-band net count rate to flux and magnitude using Swift’s effective area (corrected for UVOT’s ongoing sensitivity decline; A.
+The UVW1 stacked images showed a larger offset (approximately 5 arcsec) than the V images (approximately 1 arcsec), suggesting greater uncertainty in event mode.
+We applied a 2D Gaussian fit to the surface to locate the nucleus position for further analyses.
+Fig.
+1 shows the resulting stacked UVW1 and V-band images after applying offset corrections.
+All images show a clear detection of 3I/ATLAS.
+In the V-band images, the comet appears elongated along the direction of motion, as expected.
+The UVW1 images, free from smearing, show a spatially extended source.
+We acquired UVW1 photometry using a circular aperture with a radius of 10 arcsec, corresponding to 20,430 km and 19,084 km for the July 31 and August 19 visits, respectively.
+An elongated 10-arcsec aperture was adopted for the V-band images to account for the motion-induced elongation of the comet.
+To estimate the background level and its uncertainty, we randomly selected 2,000 points within an annulus 60–90 arcsec from the nucleus in each image.
+Around each point, we defined a region matching the size and shape of the source aperture and measured its total brightness within the region.
+The mean of these measurements was taken as the total background within the source aperture, and their standard deviation as the 1- uncertainty of the background measurement.
+1
+We also derived background-subtracted images, from which we determined the surface brightness profiles of the coma.
+For the UVW1 images, the profiles were measured using a series of concentric annuli with a width of 1 arcsec, taking the mean pixel brightness within each annulus.
+To mitigate motion blur in the V-band images, we measured the profiles from an unaffected semicircular region confined to the sunward end.
+The resulting profiles, shown in Fig.
+2, are normalized to the brightness at the nucleus of each profile.
+The comet is clearly extended beyond the PSF in each image.
+In addition, in both epochs, the UVW1 profile is broader than that of the V band, confirming the presence of an additional gas contribution in the UVW1 filter.
+We converted the V-band net count rate to band-integrated flux and magnitude using UVOT’s effective area (corrected for its ongoing sensitivity decline; A.
 A.
 Breeveld et al.
-(2011)) and a target spectrum model.
-We adopted a linearly reddened solar spectrum to model the target spectrum with the default solar spectrum included in the sbpy package (M.
-Mommert et al., 2019), and used a radius of 15 cm to derive the UVOT collecting area.
-The calculated flux, together with the same target spectrum model and Vega spectrum (R.
+(2011)) and a target spectrum model, for which we adopted a linear reddening with the default solar spectrum included in the sbpy package (M.
+Mommert et al., 2019).
+We used a radius of 15 cm to derive the UVOT collecting area.
+The calculated flux, together with the target spectrum model and Vega spectrum (R.
 C.
 Bohlin, 2014), was used to derive the apparent Vega magnitude in the V band of UVOT.
-For the V filter, the derived V-band fluxes and magnitudes are nearly insensitive to different reddening assumptions, with result variations less than 0.4%.
+The derived V-band fluxes and magnitudes are nearly insensitive to different reddening assumptions, with result variations less than 0.4%.
 Therefore, the uncertainty introduced by the reddening assumption is negligible for the photometry.
-The measured V-band flux is erg s-1 cm-2, corresponding to a V magnitude of mag.
 The net count rate of the V image also enables an estimate of the reflected continuum contribution to the UVW1 flux, allowing us to isolate the OH (A – X) emission:
@@
 F.
-A’Hearn, 1988) appropriate for the specific geocentric velocities, with a scaling applied for the heliocentric distance.
+A’Hearn, 1988) appropriate for the specific heliocentric velocities, with a scaling applied for the heliocentric distance.
 To derive water production rates from the total number of OH molecules in the aperture, we applied the vectorial model with a random OH ejection kick of 1.05 km s-1 by taking photodissociation of H2O into account (M.
@@
 Festou, 1981).
-We adopted a H2O outflow velocity of km s-1, photodissociation lifetimes of 86 000 s for H2O and 129 000 s for OH, and a branching ratio of 0.93 for water dissociation into OH (M.
+We adopted a H2O outflow velocity of km s-1, photodissociation lifetimes of 86,000 s for H2O and 129,000 s for OH, and a branching ratio of 0.93 for water dissociation into OH (M.
 R.
@@
 We compared observed and model-predicted OH molecular numbers within the same aperture.
-The observed water production rate was then obtained by scaling the modeled water production rate with this ratio.
-3 Results and Discussion
-The UVOT-V magnitude of mag corresponds to a Johnson V-mag of .
-This results in an of 174 cm after correcting for the phase angle (M.
+The water production rates were then obtained by scaling the modeled water production rate by this ratio.
+3 Results
+The measured V magnitude was mag (V flux = erg s-1 cm-2) on July 31, and mag (V flux = erg s-1 cm-2) on August 19.
+These UVOT V magnitudes correspond to Johnson V magnitudes of and , respectively, yielding values of 202 cm and 337 cm after correcting for phase angle (M.
 F.
-Ahearn et al., 1984; D.
+A’Hearn et al., 1984; D.
 G.
@@
 N.
-Bair, 2011).
-If we apply the V-R color of 0.51 mag measured on July 23 from T.
-Santana-Ros et al.
-(2025), the is 180 cm in the R band.
-The discrepancy between our results and the 300 cm reported by T.
-Santana-Ros et al.
-(2025) could originate in temporal coma development or aperture size differences.
-Our water production rate measurement depends strongly on the assumed reddening between the effective wavelengths of the V and UVW1 filters (5437.8 Å and 3325.7 Å, respectively), as shown in the left panel of Fig.
-2.
-The error bars represent 1- stochastic uncertainties.
-A possible water detection emerges for reddening assumptions larger than , while a clear 3- detection is achieved for reddening larger than .
-We then investigated reddening measurements of 3I/ATLAS from other telescopes, presented in the middle panel of Fig.
-2.
-The x-axis shows reddening values (percent per 1000 Å), while the y-axis indicates the central wavelength of the measurement bands.
-Different marker shapes denote literature sources, with colors representing the comet’s heliocentric distance during the observations (colorbar on the far right).
-A.
-Alvarez-Candal et al.
-(2025) measured per 1000 Å across 3200–5600 Å, with a central wavelength closest to ours (4381.8 Å).
-We show this reddening measurement in the left panel with a green dashed line.
+Bair, 2011), broadly consistent with other reports of measured between July 31 - Aug.
+19 (L.
+E.
+Salazar Manzano et al., 2025; T.
+Santana-Ros et al., 2025).
+Both the number of OH molecules within the aperture and the derived water production rate are highly sensitive to the assumed reddening between the effective wavelengths of the V and UVW1 filters (5437.8 Å and 3325.7 Å, respectively).
+This dependence is shown in Fig.
+3, where 4381.8 Å in the x-axis label marks the midpoint between the two filter wavelengths.
+The error bars indicate 1- stochastic uncertainties.
+During the July 31 observation, a marginal water signal (SNR) appears only when reddening assumptions exceed 24%, and even with extreme reddening assumptions up to 50%, no statistically significant detection (SNR) was obtained.
+In contrast, the August 19 observation reveals a more robust signal: a tentative detection emerges at reddening values above 13%, and a clear detection (SNR) is achieved at 24% reddening.
+We next examined reddening measurements of 3I/ATLAS obtained with other telescopes, shown in the left panel of Fig.
+4.
+The x-axis gives reddening values (percent per 1000 Å), while the y-axis marks the central wavelengths of the respective measurement bands.
+Marker shapes identify the literature sources, and the colors indicate the comet’s heliocentric distance at the time of observation.
+R.
+Rahatgaonkar et al.
+(2025) reported a color slope of approximately % per 1000 Å between 3900 Å and 5550 Å, which remained nearly constant from July 4 to August 21.
+Using their reflectance spectra (obtained via private communication), we measured the reddening (normalized color slope) at 4381.8 Å to be approximately 29%, which also remained constant through time.
+We show this reddening measurement in Fig.
+3 with a black dotted line.
 Other reddening measurements typically range from 10% to 25%, with higher reddening values at shorter wavelengths.
 This trend of increasing reddening toward shorter wavelengths is unsurprising.
-For example, a trend of increased reddening at shorter central wavelengths is presented in D.
+For example, a general trend of increased reddening of cometary dust at shorter central wavelengths is presented in D.
 Jewitt & K.
@@
 Meech, 1986).
-This trend has two sources.
-First, reddening is defined as the reflectance spectral slope normalized by the mean reflectance, not the raw slope.
-This normalization causes wavelength-dependent variations even for linear reflectance spectra, with higher values at shorter wavelengths.
-Second, reflectance spectra are often non-linear and they steepen at shorter wavelengths, further increasing reddening values.
-For 3I/ATLAS, reflectance spectra in several studies also exhibit such steepening at short wavelengths (c.f.
-T.
-Santana-Ros et al., 2025; A.
-Alvarez-Candal et al., 2025).
-To obtain reddening values at the V and UVW1 effective wavelengths for water production rate determination, we converted the middle panel’s reddening measurements from other wavelengths to UVOT wavelengths.
+This trend likely arises from two effects.
+First, reddening is defined as the reflectance spectral slope normalized by the mean reflectance, rather than the raw slope.
+This normalization introduces wavelength-dependent variations even for linear reflectance spectra, producing higher values at shorter wavelengths.
+Second, many reflectance spectra are non-linear and typically steepen toward shorter wavelengths, further enhancing reddening values.
+Although the reflectance spectrum of 3I/ATLAS does not show such steepening (R.
+Rahatgaonkar et al., 2025), the first effect alone can account for the trend observed in the left panel.
+To compare literature reddening measures directly with our Swift analysis, we converted the reported reddening to the values at effective wavelengths of UVW1 and V (Fig.
+4).
 For each measurement, we constructed a linear reflectance spectrum that, when multiplied by the solar spectrum, produces a red dust spectrum with the reported reddening value at that measurement wavelength.
-We then forward-modeled this reddened spectrum through Swift’s UVW1 and V filters and computed the reddening at UVOT effective wavelengths.
+We then forward-modeled this reddened spectrum through the transmission curves of Swift’s UVW1 and V filters and computed the reddening at UVOT’s effective wavelengths.
 Uncertainties in the converted reddening values were derived using error propagation.
 We present the converted results in the right panel of Fig.
-2.
+4.
 The y-axis still shows the original measurement wavelengths, while the x-axis displays the reddening converted to UVOT wavelengths.
 All reddening values increased after conversion, as expected from the reddening definition.
-Nearly all converted values exceed 20%, with a mean of 30.3%.
-At the 20% lower limit, the water production rate is molecules/s (SNR = 2.1), while at the mean reddening value, it reaches with SNR = 3.8, indicating clear detection.
-Our calculations assume linear reflectance spectra.
-When considering that reflectance spectra will steepen towards shorter wavelengths, actual reddening values would be higher, strengthening the water detection.
+Nearly all converted values exceed 20%, with a mean of 29%.
 The conversion accuracy depends on the wavelength difference between the original measurement (y-axis) and UVOT’s middle effective wavelength (4381.8 Å) — smaller differences yield higher accuracy.
-The measurement by A.
-Alvarez-Candal et al.
-(2025), nearest to our UVOT filter set (3200-5600 Å), converts to 38.6%, corresponding to a definitive detection of with SNR = 5.0, which corresponds to 40 kg s-1 at a heliocentric distance of 3.51 au.
-The ratio log[Af/Q(H2O)] = -24.9 places 3I/ATLAS among the most dust-rich comets observed, comparable to a small subset that includes C/1995 O1 (Hale-Bopp), C/1991 B1 (Shoemaker-Levy), and C/2009 P1 (Garradd) (M.
+The reddening of 29% measured by R.
+Rahatgaonkar et al.
+(2025) at 4381.8 Å is theoretically the most accurate.
+Additionally, this measurement matches the mean converted value, which not only validates the assumption of a linear reflectance spectrum but also strengthens our confidence in the reddening value.
+Consequently, we use a reddening of 29% for subsequent analysis and adopt the corresponding water production rate as our measurement results.
+For the July 31 visit, the number of OH molecules within the 10-arcsec aperture is (SNR = 1.5), corresponding to a water production rate of molecules s-1, indicating a marginal detection.
+In contrast, the August 19 visit yields OH molecules within the aperture (SNR = 3.9) and a water production rate of molecules s-1, and shows a clear detection.
+A summary of these results is provided in Table 1.
+Filter imaging of comets, in particular with filters that were not specifically designed for comet observations, can be contaminated by emission features from other gases.
+Recent observations from JWST and SPHEREx discovered a CO2-dominated coma (M.
+A.
+Cordiner et al., 2025; C.
+M.
+Lisse et al., 2025).
+The ionized product of CO2, CO, has the (B–X) doublet between 2884–2896 Å which lies within the bandpass of the UVW1 filter, possibly contaminating our signal.
+However, sensitive upper limits from the X-shooter spectrum (private communication) show no detectable CO A–X emission around 3674 Å.
+This rules out a significant contribution of the CO B–X doublet to the UVW1 count rate.
+Additionally, recent measurements from R.
+Rahatgaonkar et al.
+(2025) indicate the presence of Ni emission in the same spectral region.
+Thus, a potential Ni contribution cannot be ruled out, but quantifying its fraction is highly uncertain with the current data.
+4 Discussion
+The gas-to-dust ratio, expressed as log10[A(0)f/Q(H2O)], was for both visits.
+This value is consistent with solar system comets observed at heliocentric distances of 2.9–3.5 au, which span a broad range from to but generally trend toward lower values.
+The decrease in Af-to-water ratios with increasing heliocentric distance is partly attributable to the Af parameter’s assumption of constant outflow (M.
 F.
-A’Hearn et al., 1995; T.
-Farnham et al., 1997; Y.
-R.
-Fernandez, 2000; D.
-Bodewits et al., 2014).
-We note, however, that the Af values in the M.
-F.
-A’Hearn et al.
-(1995) sample were not corrected for phase angle and OH production rates were used there, which will be lower than water production rates.
+A’Hearn et al., 1995).
 Using a sublimation model222https://pdssbn.astro.umd.edu/tools/ma-evap/index.shtml (J.
@@
 F.
-A’Hearn, 1979) and assuming constant solar elevation (as would occur for slow rotation or a Sun-facing pole), a Bond albedo of 0.05%, and unit infrared emissivity, we estimate the minimum required active area to be 19.47 km2, equivalent to a circular area with radius 2.47 km.
-Observations acquired with the Hubble Space Telescope placed an upper limit on the nucleus’ radius of 2.8 km (D.
+A’Hearn, 1979) and assuming constant solar elevation (as would occur for slow rotation or a Sun-facing pole), a Bond albedo of 5%, and unity infrared emissivity, we estimate the minimum required minimal active area to be 10.6 km2 for the first visit and 7.8 km2 for the second visit.
+Observations acquired with the Hubble Space Telescope placed an upper limit on the nucleus’s radius of 2.8 km (D.
 Jewitt et al., 2025).
-If the water is produced directly by sublimation of ice within the nucleus, it would require that 20% of the surface contributes, whereas most solar system comets have active fractions around 3-5% (M.
+If the water is produced directly by sublimation of ice within the nucleus, it would require that 11–8% of the surface contributes, whereas most solar system comets have active fractions around 3–5% (M.
 F.
@@
 Keller et al., 2015).
-Gas detections beyond 2.5 au are uncommon due to the inefficiency of water sublimation at those distances.
-Only a few comets, such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), have shown OH detections at such large distances.
+The water production rates measured by Swift on July 31 (7.4 5.0 molecules/s) are comparable but slightly higher than the JWST measurement acquired on August 6 (2.19 0.08 molecules/s).
+This discrepancy may reflect systematic uncertainties if the true H2O and OH distributions deviate from the assumptions of the adopted coma models.
+Under such circumstances, using different aperture sizes — 3 ″for JWST and 10 ″for Swift—can yield different water production rates.
+For example, an extended source located outside the JWST aperture but within the Swift aperture could contribute additional OH, leading to a higher inferred production rates from the Swift measurements.
+The recent direct detection of water in 3I/ATLAS by JWST similarly shows a growing water production curve with nucleocentric distance, underscoring that nuclear sublimation alone cannot explain observed water activity and distributions at these heliocentric distances.
+Only a few comets, such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), have shown OH detections at such large distances outside of 2.5 au.
 OH emission from Bowell was detected at 5.3 au (M.
 F.
-Ahearn et al., 1984), requiring an implausibly large active area (1.6 107 km2) unless icy grains contributed significantly to water release.
+A’Hearn et al., 1984), requiring an implausibly large active area (1.6 107 km2) unless icy grains contributed significantly to water release.
 The OH production showed non-monotonic evolution, suggesting transient grain-driven activity.
@@
 Yang et al., 2025).
-If these grains are the primary source of the water and OH observed in the coma, they may explain why OH (A – X) emission was detected before any CN emission.
+If these grains are the primary source of the water and OH observed in the coma, they may explain why OH (A – X) emission was detected nearly simultaneously with CN emission.
 Typically, CN is among the most distant fragment species detected in comets due to the high fluorescence efficiency of its B – X (0,0) violet system near 3883 Å, which also benefits from lower atmospheric extinction and higher detector sensitivity compared to the OH emission near 3085 Å (c.f.
@@
 Keller et al., 2017).
-The physical processes that govern distant cometary activity and the retention or release of volatiles in such regimes remain poorly understood.
-Observational biases, limited statistics, and the challenge of detecting tenuous outgassing hinder progress.
-However, the unusual characteristics of 3I/ATLAS including early OH but no CN detection, strong coma reddening, and inferred large grains, highlight the need for long-term, multiwavelength monitoring.
-Simultaneous observations of gas and ice (e.g., via JWST) will be essential to disentangle the roles of native nucleus sublimation and extended grain activity, as well as probe the abundances of any hypervolatiles like CO and CO2.
-As noted by both M.
-J.
-Hopkins et al.
-(2025) and A.
-G.
-Taylor & D.
-Z.
-Seligman (2025), 3I/ATLAS may be water-rich, consistent with planetesimals formed in the outer regions of low-metallicity planetary systems.
-Our results offer one of the first direct tests of that hypothesis, finding indirect evidence for a large H2O production rate, but it remains to be seen if 3I/ALTAS’ coma is dominated by the other volatiles.
-The presence of substantial OH, and the inferred production of H2O, can then be used in direct comparison to 2I/Borisov.
-Borisov had a much lower intercept speed with our solar system and likely came from a more recently formed system; the high CO/H2O ratio post-perihelion was seen as evidence of formation near the CO ice line, likely of an M-type star (D.
-Bodewits et al., 2020).
-3I/ATLAS, which is moving nearly twice as fast as Borisov, is also likely to show asymmetric activity around perihelion as the thermal wave penetrates into deeper reservoirs.
-The post-perihelion behavior of 3I will expose more pristine materials, unaltered from the inbound journey.
-3I/ATLAS has a H2O production rate at 3.5 au similar to the maximum value measured for 2I/Borisov at 2.4 au (J.
-Crovisier et al., 2019; Z.
-Xing et al., 2020), and therefore we propose the following tests of the H2O-rich hypothesis:
-1.
-If 3I/ATLAS is H2O-rich, indicating an older age and formation in a low-metallicity system, then the H2O production rate will peak near perihelion, and only trace amounts of high-metallicity volatiles like CO,CO2, CN and hydrocarbon species will be found.
-2.
-If 3I/ATLAS is hyper-volatile rich (CO/CO2), indicating that its dynamic age does not necessarily represent the metallicity of where it formed, the H2O production rate will drop precipitously postperihelion as more volatile species are sublimated from depth.
-This would be similar to the CO/H2O observations of 2I/Borisov (D.
-Bodewits et al., 2020).
-If the first hypothesis is born out, and 3I/ATLAS is indeed from an older low-metallicity system, than this may suggest that there is a part of the planetesimal formation or removal process that may have ejected more mass from forming systems in order to properly place the large size of 3I into context (A.
-G.
-Taylor & D.
-Z.
-Seligman, 2025).
-Given the ISO production of low-metallicity systems would likely come from planetary interactions (S.
-N.
-Raymond et al., 2018) or stripping by stellar flybys (S.
-Pfalzner et al., 2021) an overabundance of large, H2O-rich ISO’s could serve as a key indicator of ejection conditions.
+With the recent spectral confirmation of CN in the coma of 3I/ATLAS we can derive log10(CN/OH) values to compare to typical solar system comet values.
+Reported CN production rates between August 10 and 19 span a wide range, from 0.2 to 7.81024 molecules/s (D.
+Schleicher, 2025; R.
+Rahatgaonkar et al., 2025; L.
+E.
+Salazar Manzano et al., 2025).
+To derive , we obtained the OH production rate () by multiplying the measured water production rate by the branching ratio (0.93) and interpolating to each , then computed for all CN measurements, yielding a median value of about , which is resilient to outliers.
+This would make 3I/ATLAS depleted in CN compared to even depleted solar system comets, which exhibit log10(CN/OH)=-2.69 (M.
+F.
+A’Hearn et al., 1995).
+However, this estimate is highly model-dependent, since is derived from water production rates using vectorial model assumptions that may not hold at large heliocentric distances.
+The potential role of icy grains highlights the complexity of the physical processes governing distant activity and volatile retention or release, which remain poorly constrained due to observational biases, limited statistics, and the difficulty of detecting tenuous outgassing.
+Therefore, the assumptions of standard vectorial models — steady-state production, symmetric outflow, and nucleus-centered sources — may break down, leading to systematic uncertainty in the derived water and OH production rates.
+In addition, our preliminary radial profiles indicate a more concentrated OH distribution than predicted by the vectorial model, implying that both and may be overestimated.
+Consequently, the aperture-integrated number of OH molecules provides a more direct and less model-dependent constraint, while the conversion to production rates should be treated with caution.
+The unusual characteristics of 3I/ATLAS including early OH, a depleted and possibly varying CN/H2O ratio (D.
+Schleicher, 2025; R.
+Rahatgaonkar et al., 2025; L.
+E.
+Salazar Manzano et al., 2025), strong coma reddening, and inferred large grains, highlight the need for long-term, multiwavelength monitoring.
+Simultaneous observations of gas and ice will be essential to disentangle the roles of native nucleus sublimation and extended grain activity.
 5 Conclusion
@@
 This confirms the presence of water in the third known interstellar object.
-The derived production rate of molecules s-1 (40 kg s-1) is notable given the object’s large heliocentric distance of 3.51 au, where water sublimation is typically inefficient.
+Based on a dust reddening of 29% at 4381.8Å, we obtained only a marginal detection of OH emission during the July 31–August 1 visit at au, corresponding to a water production rate of ( molecules s-1.
+However, subsequent observations on August 18–20 revealed a clear detection, corresponding to a water production rate of molecules s-1 (40 kg s-1) at au.
+This is notable given that water sublimation is typically inefficient at such large heliocentric distances.
 This level of activity is consistent with a reasonably-sized active area assuming equilibrium sublimation, but could also result from the presence of large icy grains in the coma, as suggested by contemporaneous near-infrared observations.
-If the coma of 3I/ATLAS remains H2O dominated and depleted in higher-metallicity volatiles like CO, CN, and CO2, the relatively large size of the interstellar object may be a telling clue about the planetesimal and mass loss processes of early low-metallicity systems.
-The detection of OH emission prior to any detection of CN is unusual and may indicate differences in grain composition or volatile distribution compared to solar system comets (C.
-Opitom et al., 2025; A.
-Alvarez-Candal et al., 2025).
+If the coma of 3I/ATLAS becomes H2O-dominated during and after perihelion and shows lower abundances of higher-metallicity volatiles like CO, CN, and CO2, a trend that would be the inverse of 2I/Borisov, this may support the early and low-metallicity formation hypothesis, and the derived large size of the nucleus could be indicative of a key knowledge gap in low-metallicity system planetesimal formation and loss mechanisms.
+The detection of OH emission nearly simultaneously to the detection of CN is unusual and may indicate differences in grain composition or volatile distribution compared to solar system comets (C.
+Opitom et al., 2025).
 While similar behavior has been observed in comets such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), the processes governing distant activity and the release of volatiles remain poorly constrained.
@@
 This detection demonstrates the critical role of UV space-based observations in tracking the onset and evolution of activity in distant or dynamically unusual objects.
+The relatively weak CN abundances relative to H2O and the apparent large abundance of CO2 relative to H2O would seem to contradict the theorized low-metallicity formation.
+The upcoming perihelion passage will likely expose less-altered materials as depleted layers are stripped from its surface, and we can continue to test the low-metallicity hypothesis.
 Continued monitoring of 3I/ATLAS will help clarify the thermal and compositional properties of interstellar ices and improve our understanding of volatile retention and release in planetary systems beyond our own.